The reason there is danger in exposure to high energy radiation is that the radiation can rupture chemical bonds. In some cases cancer can be caused by it. A C-C single bond has a bond energy of aprox 348 kJ/mol. What wavelength ( in nm) of light is required to provide sufficient energy to break the C-C bond?
Answer #1

The energy of C-C single bond is given by 348 kJ/mol = 348000 J/mol

Bond energy of one C-C single bond is 348000/(6.022 x 1023) J = 5.779 x 10-19 J

Now we have E = hc/λ

Where E = 5.779 x 10-19 J, h = 6.626 x 10-34 Js, c = 3 x 108 m/s

λ = hc/E = (6.626 x 10-34 Js x 3 x 108 m/s)/( 5.779 x 10-19 J) = 3.439 x 10-7 m = 343.9 nm

Hence the wavelength of light required is 343.9 nm.